If x is an odd integer, then x


2 + 3x + 5 is odd. (Use Direct Proof)


1
Expert's answer
2022-04-26T00:18:35-0400

Any odd number can be presented as: "x=2n+1," where "n" is an integer. Thus, we get: "2+3(2n+1)+5=6n+10=2(3n+5)". As we can see, the latter number is even. Thus, the statement is wrong.